# Pentagram maps over rings, Grassmannians, and skewers

### v2

@@ 1 Introduction
The pentagram map has been generalized in several different directions. One can intersect diagonals joining vertices that are further apart, consider polygons in higher-dimensional spaces, etc. [10, 11, 6]. Here, we are interested in other kinds of generalizations, which are based on replacing conventional notions of points and lines by more general objects with similar properties. One example of such a map is the Grassmann pentagram map, defined in [4] and further studied in [14, 2]. In this setting, the role of points is played by elements of $\mathrm{Gr}(k,3k)$, i.e. $k$-dimensional subspaces in a $3k$-dimensional vector space, while lines are elements of $\mathrm{Gr}(2k,3k)$. Any two generic points $V_{1},V_{2}\in\mathrm{Gr}(k,3k)$ belong to a single line $V_{1}\oplus V_{2}\in\mathrm{Gr}(2k,3k)$, and any two generic lines $W_{1},W_{2}\in\mathrm{Gr}(2k,3k)$ meet at a point $W_{1}\cap W_{2}\in\mathrm{Gr}(k,3k)$. The Grassmann pentagram map takes a (cyclically labeled) sequence of points $V_{i}\in\mathrm{Gr}(k,3k)$, which we think of as a polygon in the Grassmannian $\mathrm{Gr}(k,3k)$, to a new sequence given by $(V_{i}\oplus V_{i+2})\cap(V_{i+1}\oplus V_{i+3})$. In other words, just like the classical pentagram map, the Grassmann map is defined by intersecting diagonals connecting second-nearest vertices. Furthermore, the classical pentagram map is nothing else but the $k=1$ version of the Grassmann map.
@@
In the present paper, we develop a common framework for these types of pentagram-like maps. Specifically, we show that such maps can be viewed as pentagram maps in the projective plane over an appropriate ring. In general, those rings need not be division rings or commutative. Specifically, we show that the Grassmann pentagram map corresponds to the ring of matrices, while the skewer map is the pentagram map over the ring $\mathbb{R}[\varepsilon]/(\varepsilon^{2})$ of dual numbers.
Furthermore, we prove that the pentagram map is integrable for any stably finite ground ring $R$, thus generalizing the results of [4, 14] on integrability of the Grassmann map and answering the question of [21] regarding integrability of the skewer map. Integrability is understood in this paper as the existence of a Lax form. A discrete dynamical system $\phi$ on a set $X$ (i.e., a map $\phi\colon X\to X$) is said to admit a Lax form if to every $x\in X$ one can assign an operator in such a way that the operators assigned to $x$ and $\phi(x)$ are similar. In the setting of pentagram maps over a ring $R$, these operators are pseudo-difference operators with coefficients in $R$. When $R$ is a field, we recover the Lax form of the classical pentagram map given in [6, 8].
The existence of a Lax form is one of the characteristic features of integrable dynamical systems. In particular, for the classical pentagram map (i.e., for the pentagram map over a field), the Lax form has been a central tool for establishing the map’s dynamical properties, such as quasi-periodicity, cf. [18, 25]. The study of dynamical aspects of integrability for pentagram maps over rings is beyond the scope of the present paper. However, we do expect that quasi-periodicity-type results can be extended from fields to a fairly large class of rings, in particular to finite-dimensional algebras.
The structure of the paper is as follows.
In Section 2, we recall the definition and standard properties of stably finite rings. Details can be found, e.g., in [13, Chapter 1, §1B].
@@
Sections 3.2 and 3.4 are devoted to projective spaces over the full matrix ring, which is the first of our two main examples. The second main example, the ring of dual numbers, is discussed in Section 3.5.
@@
In Section 5, we introduce our main technical tool: difference operators with coefficients in a ring. After discussing basic properties of such operators, we use that language to give an algebraic description of the moduli space of polygons in the projective space over a ring. The main result of the section, Theorem 5.7, is a ring counterpart of [8, Proposition 3.3] which gives a description of real polygons in terms of difference operators. We note that the proof in the ring case is substantially more technical due to the lack of commutativity and essentially relies on the stable finiteness condition.
In Section 6, we establish integrability of the pentagram over an arbitrary stably finite ring. As a first step, in Section 6.1, we reformulate the pentagram map in the language of difference operators. The main result of the section is Theorem 6.5, which generalizes [8, Theorem 1.1] and essentially says that, in terms of difference operators, the pentagram map is described by a linear equation. This is used in Section 6.3 to give a Lax formulation of the pentagram map. It is given in terms of pseudo-difference operators, which are introduced in Section 6.2.
Finally, in Section 7, we use the Lax form to give invariants (i.e., conserved quantities, or first integrals) of the ring pentagram map on closed polygons. In the case of a field, these invariants are known to form a maximal Poisson-commuting family [15]. We expect that similar results can be obtained for much bigger classes of rings.
Acknowledgments. The authors are grateful to Boris Khesin, Nicholas Ovenhouse, Richard Schwartz, and Max Weinreich for fruitful discussions. This work was supported by NSF grant DMS-2008021.
@@ 2 Stably finite rings
In this section, we recall the definition and basic properties of stably finite rings – the main class of rings relevant to the present paper; for details see [13, Chapter 1, §1B]. In what follows, all rings are assumed to be unital and non-trivial (i.e. have a multiplicative identity element $1\neq 0$), but may be non-commutative and/or have zero divisors. All modules are supposed to be unitary.
@@
For projective (in particular, free) modules, this condition can also be reformulated as follows.
@@
Suppose $R$ is a ring, and $M$ is a projective $R$-module. Then the following conditions are equivalent.
@@
    the ring $\mathrm{End}_{R}(M)$ is Dedekind-finite (recall that a ring is Dedekind-finite if $ab=1$ implies $ba=1$).
@@
    if $M\simeq M\oplus N$ for some other $R$-module $N$, then $N=0$.
Proof.
@@
$3\Rightarrow 1$. Suppose that $M\simeq M\oplus N$ implies $N=0$. Let $\phi\in\mathrm{End}_{R}(M)$ be surjective. Then, since $M$ is projective, the exact sequence $0\to\mathrm{Ker}\,\phi\to M\to M\to 0$ splits, and $M\simeq M\oplus\mathrm{Ker}\,\phi$. So, $\mathrm{Ker}\,\phi=0$, proving that $M$ is Hopfian. ∎
@@
One of the most important properties of stably finite rings is that they have an invariant basis number property:
@@
So, for every finite rank free module over a stably finite ring, the notion of rank is well-defined.
@@
The first claim follows from Proposition 2.2, combined with the isomorphism $\mathrm{End}_{R}(M)\simeq\text{M}_{n}(R)$ for a free rank $n$ right module $M$. The second claim follows from the identification $\text{M}_{1}(R)\simeq R$. ∎
The following says that all rings that are “not too large” are stably finite.
@@
All Noetherian rings are stably finite.
@@
Let $M$ be a finite rank free module over a Noetherian ring $R$. Then $M$ is a Noetherian module. Let $\phi\in\mathrm{End}_{R}(M)$ be surjective. By the Noetherian property, the chain of submodules $\mathrm{Ker}\,\phi\subset\mathrm{Ker}\,\phi^{2}\subset\dots$ must stabilize: $\mathrm{Ker}\,\phi^{n}=\mathrm{Ker}\,\phi^{n+1}$ for some $n$. Suppose $v\in\mathrm{Ker}\,\phi$. By surjectivity, there is $w\in M$ such that $v=\phi^{n}(w)$. Then $\phi^{n+1}(w)=\phi(v)=0$, so $w\in\mathrm{Ker}\,\phi^{n+1}$. But $\mathrm{Ker}\,\phi^{n}=\mathrm{Ker}\,\phi^{n+1}$, so $v=\phi^{n}(w)=0$. Thus, $\phi$ is an isomorphism. ∎
As a corollary, all Artinian rings are stably finite. This can also be proved directly, as follows. Assume $R$ is Artinian, and let $M$ be a free finite rank module over $R$. Then $M$ is also Artinian and hence can not have proper submodules isomorphic to $M$ (otherwise, repetition gives an infinite descending chain of submodules). So, $M\simeq M\oplus N$ implies $N=0$, meaning that $R$ is stably finite.
@@ 3 Projective spaces over rings
This section is an introduction into projective geometry over rings. We note that while ring geometry is a well-studied subject (see [24]), the setting of stably finite rings appears to be new.
@@ 3.1 The projectivization of a module
There exist several, generally inequivalent, ways to define projective spaces over rings. Here we adapt the definition given in [23]. Although we will mainly be interested in projective spaces obtained from free modules, we first define the projectivization of an arbitrary module.
@@
Let $M$ be a (left or right) $R$-module. A subspace $N\subset M$ is a free submodule which admits a direct complement in $M$ (i.e., there is a submodule $K\subset M$, not necessarily free, such that $M=N\oplus K$).
@@
One can projectivize both left modules and right modules.
@@
Given $R$-modules $M,N$ of the same kind (both left or both right), a projective transformation $\mathbb{P}(M)\to\mathbb{P}(N)$ is a map induced by a module isomorphism $M\to N$.
For a module $M$, its projectivization $\mathbb{P}(M)$ may also be defined as the set of unimodular elements modulo scalars. For any module $M$, let $M^{*}=\mathrm{Hom}_{R}(M,R)$ be its dual module.
@@
Let $M$ be an $R$-module. An element $v\in M$ is called unimodular if there exists $\xi\in M^{*}$ such that $\xi(v)=1$ (equivalently, if there exists $\xi\in M^{*}$ such that $\xi(v)$ is a unit).
@@
We give a proof for the case of left modules (the case of right modules is analogous). Suppose $P\subset M$ is a rank one subspace. Then, by definition, $P$ is freely generated over $R$ by some $v\in M$, and is a direct summand in $M$, i.e. $M=P\oplus N$. Therefore, for each $w\in M$ there exists unique $r\in R$ and $u\in N$ such that $w=rv+u$. Define a homomorphism $\xi\colon M\to R$ by $\xi(rv+u):=r$. Then $\xi(v)=1$, ensuring $v$ is unimodular.
@@
We give a proof for the case of left modules (the case of right modules is analogous). Clearly, if $r$ is a unit, then $Rv=Rrv$. Conversely, assume that $Rv=Rw$. Then $w\in Rv$, so $w=rv$ for some $r\in R$. Likewise, $v=r^{\prime}w$ for some $r^{\prime}\in R$. So, $w=rv=rr^{\prime}w$. Since $w$ is unimodular, this implies $rr^{\prime}=1$. Similarly, $r^{\prime}r=1$. So, $r$ is a unit. ∎

### v3

@@ 1.1 Overview
The pentagram map has been generalized in several different directions. One can intersect diagonals joining vertices that are further apart, consider polygons in higher-dimensional spaces, etc. [7, 11, 12]. Here, we are interested in other kinds of generalizations which are based on replacing conventional notions of points and lines by more general objects with similar properties. One example of such a map is the Grassmann pentagram map, defined in [5] and further studied in [2, 17]. In this setting, the role of points is played by elements of $\mathrm{Gr}(k,3k)$, i.e. $k$-dimensional subspaces in a $3k$-dimensional vector space, while lines are elements of $\mathrm{Gr}(2k,3k)$. Any two generic points $V_{1},V_{2}\in\mathrm{Gr}(k,3k)$ belong to a single line $V_{1}\oplus V_{2}\in\mathrm{Gr}(2k,3k)$, and any two generic lines $W_{1},W_{2}\in\mathrm{Gr}(2k,3k)$ meet at a point $W_{1}\cap W_{2}\in\mathrm{Gr}(k,3k)$. The Grassmann pentagram map takes a (cyclically labeled) sequence of points $V_{i}\in\mathrm{Gr}(k,3k)$, which we think of as a polygon in the Grassmannian $\mathrm{Gr}(k,3k)$, to a new sequence given by $(V_{i}\oplus V_{i+2})\cap(V_{i+1}\oplus V_{i+3})$. In other words, just like the classical pentagram map, the Grassmann map is defined by intersecting diagonals connecting second-nearest vertices. Furthermore, the classical pentagram map is nothing else but the $k=1$ version of the Grassmann map.
@@
In the present paper, we develop a common framework for these types of pentagram-like maps. Specifically, we show that such maps can be viewed as pentagram maps in the projective plane over an appropriate ring. In general, those rings need not be division rings or commutative. We show that the Grassmann pentagram map corresponds to the ring of matrices, while the skewer map is the pentagram map over the ring $\mathbb{R}[\varepsilon]/(\varepsilon^{2})$ of dual numbers.
Furthermore, we prove that the pentagram map is integrable for any stably finite ground ring $R$, thus generalizing the results of [5, 17] on integrability of the Grassmann map and answering the question of [25] regarding integrability of the skewer map. The formal definition of stably finite rings will be given in Section 2, but we emphasize here that this is a large class of rings which includes, amongst others, all commutative and all Noetherian rings.
Integrability is understood in this paper as the existence of a Lax form. A discrete dynamical system is said to admit a Lax form if it can be written in the form $L\mapsto A^{-1}LA$ for certain operators $L$, AA (for details, see Section 6.3). In the setting of pentagram maps over a ring $R$, these operators are pseudo-difference operators with coefficients in $R$. When $R$ is a field, we recover the Lax form of the classical pentagram map given in [7, 9].
The existence of a Lax form is one of the characteristic features of integrable dynamical systems. In particular, for the classical pentagram map (i.e., for the pentagram map over a field), the Lax form has been a central tool for establishing the map’s dynamical properties, such as quasi-periodicity, cf. [22, 30]. The study of dynamical aspects of integrability for pentagram maps over rings is beyond the scope of the present paper. However, we do expect that quasi-periodicity-type results can be extended from fields to finite-dimensional algebras over fields.
Our main results can be summarized as follows:
Theorem 1.1.
Let $R$ be a stably finite ring. Then there is a natural one-to-one correspondence between the set of projective equivalence classes of polygons in the left projective space $\mathbb{P}^{d}(R)$ and an appropriate quotient of the space of degree $d+1$ left difference operators with coefficients in $R$. In terms of difference operators, the pentagram map in $\mathbb{P}^{2}(R)$ is described by a linear equation
$\widehat{\mathcal{D}}_{+}\mathcal{D}_{-}=\widehat{\mathcal{D}}_{-}\mathcal{D}_{+}$
where the operators $\mathcal{D}_{\pm}$ are associated to a polygon, and $\widehat{\mathcal{D}}_{\pm}$ to its image under the pentagram map. As a corollary, the pentagram map over $R$ admits a Lax representation.
We note that while similar results in the case of a field were obtained by the second author in [9], the arguments from that paper do not seem to generalize to the ring setting, and the construction of the present paper is essentially different. The main technical challenges to overcome are possible non-commutativity of the base ring along with the impossibility of resorting to general-position-type arguments.
We also note that while there are versions of the pentagram map in projective spaces of higher dimension, we chose to not pursue this direction here, as it comes with additional technical difficulties related to more complicated combinatorics of multi-dimensional pentagram maps (intersection of a collection of planes as opposed to two lines).
1.2 Structure of the paper
In Section 2, we recall the definition and standard properties of stably finite rings. Details can be found, e.g., in [15, Chapter 1, §1B]. We also include a concise glossary, Table 1, of terms from ring and module theory that are relevant to this paper.
@@ 1.2 Structure of the paper
Sections 3.2 and 3.4 are devoted to projective spaces over the full matrix ring, which is the first of our two main examples. The second main example, the ring of dual numbers, is discussed in Section 3.5. Additionally, in Section 3.6, we discuss projective spaces over path algebras of quivers, including the algebra of triangular matrices.
@@
In Section 5, we introduce our main technical tool: difference operators with coefficients in a ring. After discussing basic properties of such operators, we use that language to give an algebraic description of the moduli space of polygons in the projective space over a ring. The main result of the section, Theorem 5.7, is a ring counterpart of [9, Proposition 3.3] which gives a description of real polygons in terms of difference operators. We note that the proof in the ring case is substantially more technical due to the lack of commutativity and fundamentally relies on the stable finiteness condition.
In Section 6, we establish integrability of the pentagram over an arbitrary stably finite ring. As a first step, in Section 6.1, we reformulate the pentagram map in the language of difference operators. The main result of the section is Theorem 6.5, which generalizes [9, Theorem 1.1] and says that, in terms of difference operators, the pentagram map is described by a linear equation. This is used in Section 6.3 to give a Lax form of the pentagram map. It is given in terms of pseudo-difference operators, which are introduced in Section 6.2.
Finally, in Section 7, we use the Lax form to give invariants (i.e., conserved quantities, or first integrals) of the ring pentagram map on closed polygons. These invariants are valued in the cyclic space of $R$ which is a non-trivial Abelian group unless $R$ is a commutator ring, i.e., is spanned over $\mathbb{Z}$ by expressions of the form abbaab-ba. In particular, we get non-trivial invariants for any commutative ring (whose cyclic space is the ring itself) and the full matrix ring (whose cyclic space coincides with the ground field). In the case when the ground ring is $\mathbb{R}$, these invariants (in a slightly more general setting of twisted polygons) are known to form a maximal Poisson-commuting family [18]. We expect that similar results can be obtained in the ring setting.
Acknowledgments. The authors are grateful to Boris Khesin, Nicholas Ovenhouse, Richard Schwartz, Max Weinreich, and anonymous referees for useful comments and fruitful discussions. This work was supported by NSF grant DMS-2008021 and the Simons Foundation through its Travel Support for Mathematicians program. A part of this work was done during A.I.’s visit to Max Planck Institute for Mathematics in Bonn. A.I. would like to thank the Institute’s faculty and staff for their support and stimulating atmosphere.
@@ 2 Stably finite rings
Unital ring Ring with multiplicative identity.
Opposite ring Ring with same elements and addition but reverse multiplication order.
Dedekind-finite ring Ring for which $ab=1\implies ba=1$.
Left Noetherian ring Ring for which any ascending chain of left ideals must stabilize.
Cyclic space of a ring Quotient $R/[R,R]$, where $[R,R]$ is the $\mathbb{Z}$-span of commutators.
Commutator ring Ring $R$ such that $[R,R]=R$.
Endomorphism Homomorphism from a module to itself.
Split exact sequence Short exact sequence isomorphic to $0\to A\to A\oplus B\to B\to 0$.
Unital module Module $M$ over unital ring such that $1v=v$ for all $v\in M$.
Cyclic module Module generated by a single element.
Free module Module that has a basis (need not be finite in length).
Finite-rank free module Free module that admits a basis of finite length.
Hopfian module Module for which any surjective endomorphism is also an isomorphism.
Noetherian module Module for which any ascending chain of submodules must stabilize.
Stably finite ring Ring for which any finite-rank free module is Hopfian.
Ring with IBN property Ring for which all bases of a finite-rank free module have same length.
Rank of a free module Number of elements in a basis. Well-defined for rings with IBN property.
Table 1: Glossary of terms related to rings and modules
In this section, we recall the definition and basic properties of stably finite rings – the main class of rings relevant to the present paper; for details, see [15, Chapter 1, §1B]. One of the main results of the section is that such rings have an invariant basis number property, which means that any finitely generated free module over such a ring has a well-defined rank, just like any vector space over a field has well-defined dimension.
For the reader’s convenience, we include a concise table of relevant terms from ring theory that will be used frequently throughout the paper in Table 1. More detailed definitions and examples are given throughout the paper. In what follows, all rings are assumed to be unital and non-trivial (i.e. have a multiplicative identity element $1\neq 0$), but may be non-commutative and/or have zero divisors. All modules are supposed to be unitary.
@@
Example 2.2
A finite-dimensional vector space is a Hopfian module. The vector space $\mathbb{R}[x]$ of single-variable polynomials is not Hopfian, since the differentiation map is surjective but not an isomorphism.
For free modules, the Hopfian condition can also be reformulated using the notion of a Dedekind-finite ring. Recall that a ring is Dedekind-finite if $ab=1$ implies $ba=1$.
Example 2.3
A common example of a non-Dedekind-finite ring is the ring of endomorphisms of the vector space $\mathbb{R}^{\mathbb{N}}$ of real-valued sequences $x_{1},x_{2},\dots\in\mathbb{R}$. Consider left shift and right shift operators on $\mathbb{R}^{\mathbb{N}}$, defined by $L(x_{1},x_{2},\dots)=x_{2},x_{3},\dots$, $R(x_{1},x_{2},\dots)=0,x_{1},x_{2},\dots$. Then $L\circ R$ is the identity, but $R\circ L$ is not.
@@
Suppose $R$ is a ring, and $M$ is a free $R$-module. Then the following conditions are equivalent.
@@
    the ring $\mathrm{End}_{R}(M)$ is Dedekind-finite
@@
    if $M\simeq M\oplus N$ for some other $R$-module NN, then $N=0$.
In what follows, the $\oplus$ symbol might mean external or internal direct sum, depending on the context. Here, $M\simeq M\oplus N$ means that $M$ is isomorphic to the external direct sum of $M$ and $N$, since an internal direct sum would not make sense in this setting. On the other hand, writing $M=N\oplus O$ for a given module $M$ would mean that $M$ is the internal direct sum of its submodules $N$ and $O$.
Proof of Proposition 2.4.
@@
$3\Rightarrow 1$. Suppose that $M\simeq M\oplus N$ implies $N=0$. Let $\phi\in\mathrm{End}_{R}(M)$ be surjective. Then, since $M$ is free, the exact sequence $0\to\mathrm{Ker}\,\phi\to M\to M\to 0$ splits, and $M\simeq M\oplus\mathrm{Ker}\,\phi$. So, $\mathrm{Ker}\,\phi=0$, proving that $M$ is Hopfian. ∎
@@
One of the most important properties of stably finite rings is that they have an invariant basis number (IBN) property:
@@
Thus, for every finite rank free module over a stably finite ring, the notion of rank is well-defined.
@@
The first claim follows from Proposition 2.4, combined with the isomorphism $\mathrm{End}_{R}(M)\simeq\mathrm{M}_{n}(R)$ for a free rank $n$ right module $M$. The second claim follows from the identification $\mathrm{M}_{1}(R)\simeq R$. ∎
Recall that a ring is left Noetherian if any ascending chain $I_{1}\subset I_{2}\subset\dots$ of left ideals eventually stabilizes, i.e., $I_{k}=I_{k+1}=\dots$ for some $k$. Most rings encountered in applications are Noetherian.
Example 2.10
Any finitely generated algebra over a field is a Noetherian ring. On the other hand, the ring $\mathbb{R}[x_{1},\dots]$ of polynomials in infinitely many variables is not Noetherian, since we have a strictly ascending chain of ideals $\langle x_{1}\rangle\subset\langle x_{1},x_{2}\rangle\subset\dots$
@@
All left Noetherian rings are stably finite.
@@
Let $M$ be a finite rank free left module over a left Noetherian ring $R$. Then $M$ is a Noetherian module. Let $\phi\in\mathrm{End}_{R}(M)$ be surjective. By the Noetherian property, the chain of submodules $\mathrm{Ker}\,\phi\subset\mathrm{Ker}\,\phi^{2}\subset\dots$ must stabilize: $\mathrm{Ker}\,\phi^{n}=\mathrm{Ker}\,\phi^{n+1}$ for some $n$. Suppose $v\in\mathrm{Ker}\,\phi$. By surjectivity, there is $w\in M$ such that $v=\phi^{n}(w)$. Then $\phi^{n+1}(w)=\phi(v)=0$, so $w\in\mathrm{Ker}\,\phi^{n+1}$. But $\mathrm{Ker}\,\phi^{n}=\mathrm{Ker}\,\phi^{n+1}$, so $v=\phi^{n}(w)=0$. Thus, $\phi$ is an isomorphism.
Likewise, all right Noetherian rings are stably finite.
@@ 3 Projective spaces over rings
This section is an introduction into projective geometry over rings. We note that while ring geometry is a well-studied subject, the setting of stably finite rings considered here appears to be more general than what can be found in the existing literature. For instance, [28, 29] only consider rings of stable rank $2$. Any such ring is stably finite, but not every stable finite ring has stable rank $2$. For instance, the ring of integers, which is commutative and hence stably finite, has stable rank $3$. Furthermore, there exist commutative, and hence stably finite, rings of arbitrary finite and even infinite stable rank [16, 27].
@@ 3.1 The projectivization of a module
There exist several, generally inequivalent, ways to define projective spaces over rings. Here we adapt the definition given in [28, Section 2]. Although we will mainly be interested in projective spaces obtained from free modules, we first define the projectivization of an arbitrary module. We only consider the case of left modules. Right modules are considered analogously.
@@
Let $M$ be an $R$-module. A subspace $N\subset M$ is a free submodule which admits a direct complement in $M$ (i.e., there is a submodule $K\subset M$, not necessarily free, such that $M=N\oplus K$).
Example 3.2
Consider $\mathbb{Z}$ as a module over itself. $2\mathbb{Z}\subset\mathbb{Z}$ is its free submodule but not a subspace, since it does not admit a complement.
@@
Given $R$-modules $M,N$, a projective transformation $\mathbb{P}(M)\to\mathbb{P}(N)$ is a map induced by a module isomorphism $M\to N$.
For a module $M$, its projectivization $\mathbb{P}(M)$ may also be defined as the set of unimodular elements modulo scalars.
@@
Let $M$ be an $R$-module, and let $M^{*}=\mathrm{Hom}_{R}(M,R)$ be its dual module. An element $v\in M$ is called unimodular if there exists $\xi\in M^{*}$ such that $\xi(v)=1$ (equivalently, if there exists $\xi\in M^{*}$ such that $\xi(v)$ is a unit).
Example 3.7
In the $\mathbb{Z}$-module $\mathbb{Z}^{n}$, an element $(k_{1},\dots,k_{n})$ is unimodular if and only if $\mathrm{gcd}(k_{1},\dots,k_{n})=1$.
@@
Suppose $P\subset M$ is a rank one subspace. Then, by definition, $P$ is freely generated over $R$ by some $v\in M$, and is a direct summand in $M$, i.e. $M=P\oplus N$. Therefore, for each $w\in M$ there exists unique $r\in R$ and $u\in N$ such that $w=rv+u$. Define a homomorphism $\xi\colon M\to R$ by $\xi(rv+u):=r$. Then $\xi(v)=1$, ensuring $v$ is unimodular.
@@
Clearly, if $r$ is a unit, then $Rv=Rrv$. Conversely, assume that $Rv=Rw$. Then $w\in Rv$, so $w=rv$ for some $r\in R$. Likewise, $v=r^{\prime}w$ for some $r^{\prime}\in R$. So, $w=rv=rr^{\prime}w$. Since $w$ is unimodular, there is $\xi\in M^{*}$ such that $\xi(w)=1$. So,
$rr^{\prime}=rr^{\prime}\xi(w)\stackrel{{\scriptstyle(*)}}{{=}}\xi(rr^{\prime}w)=\xi(w)=1,$
where ()(*) holds because $\xi\colon M\to R$ is a homomorphism of left $R$-modules. Similarly, $r^{\prime}r=1$. So, $r$ is a unit. ∎